# A Methodology for Incompleteness-Tolerant and Modular Gradual Semantics for Argumentative Statement Graphs

### v5

@@ 1 Introduction
Argumentation frameworks have emerged as powerful tools for reasoning about and resolving conflicting information in complex environments (see Baroni et al. (2018a); Atkinson et al. (2017) for overviews). In recent years, gradual semantics (GS) have shown great promise in extending these frameworks, particularly in the realm of quantitative bipolar argumentation frameworks (QBAFs) Baroni et al. (2018b), where arguments have weights and may be related by attack and support relations. These approaches thus allow for both negative and positive, respectively, influences between arguments to be accounted for, arguably aligning more closely with human judgment Polberg and Hunter (2018). GS have found applications in diverse (and often human-centric) areas, from judgmental forecasting Irwin et al. (2022) to explainable AI Vassiliades et al. (2021); Cyras et al. (2021), by allowing for quantitative, and thus in some cases more nuanced, evaluations of arguments beyond the traditional accept/reject dichotomy in extension-based semantics Dung (1995); Cayrol and Lagasquie-Schiex (2005). These quantitative evaluations, typically referred to as strengths, can be particularly useful in cases where the information in argumentation frameworks is incomplete, since strengths can account for this uncertainty quantitatively.
@@
However, most of the existing work on GS (see Baroni et al. (2019) for an overview) has focused on formalisms based on abstract arguments Dung (1995), treated as atomic entities without internal structure. This abstraction, while useful in many contexts, fails to capture the rich logical structure that is often present in real-world arguments. Meanwhile, various forms of structured argumentation frameworks (see Besnard et al. (2014) for an overview) allow to represent the internal composition of arguments and the relationships between their components, and thus offer a more expressive and realistic model in agent-based domains, such as user modeling  Hadoux et al. (2023), scientific debates Cramer and Dauphin (2020), and model reconciliation settings Vasileiou et al. (2024). However, this greater reliance on the logical (e.g., background or contextual) knowledge can lead to more problematic cases where this information is incomplete. Yet, it is only recently that the study of GS for structured argumentation has received attention, e.g., for deductive argumentation Besnard and Hunter (2014) by Heyninck et al. Heyninck et al. (2023), for assumption-based argumentation (ABA) Cyras et al. (2017) by Skiba et al. Skiba et al. (2023), for ASPIC+ Modgil and Prakken (2014) by Prakken Prakken (2024) and for a restricted type of structured argumentation in the form of statement graphs (SGs) Hecham et al. (2018) by Jedwabny et al. Jedwabny et al. (2020).
@@
Figure 1 illustrates a debate about climate change as an SG, which uses a restricted type of structured arguments in the form of statements, each consisting of a premise (a compound of literals, possibly $\top$ true) and a claim (a literal). A typical statement ($\alpha_{1}$) may combine evidence about rising CO2 emissions ($a$) and increasing global temperatures ($b$), given that they are both supported by facts (see $\alpha_{2}$ and $\alpha_{3}$), to support the claim that climate change is human-caused ($c$). The GS proposed in Jedwabny et al. (2020) uses T-norms and T-conorms from fuzzy logic to aggregate the strengths of a statement’s components, but this GS faces challenges when dealing with incomplete information since it relies on an assumption of complete knowledge and presence of a statement’s supporting elements. This requirement can limit its applicability in real-world scenarios, where such information is almost always incomplete. For example, in Figure 1, the strength of the statement $\alpha_{1}$ is reduced (e.g., in the eyes of a modelled audience) by a statement ($\alpha_{4}$) questioning the reliability of CO2 measurements, without being evidenced (there is no support for $d$). To date, existing approaches to GS struggle to handle such incompleteness, potentially leading to oversimplified or incorrect conclusions, and thus poor modelling of debates.
To address this limitation, we introduce a novel methodology for GS in SGs that can naturally accommodate complete and incomplete information. Our approach is modular in that it allows for different GS to be instantiated based on the requirements of a given task and context. Specifically, our GS can be instantiated with any existing GS for QBAFs, which is used to dialectically evaluate the literals in a premise, based on their attacking and supporting evidence, before these “intermediate” evaluations are aggregated, depending on the statement’s construction, to give a final strength for the statement. Thus, our methodology combines some of the expressiveness of structured argumentation with the flexibility of GS for QBAFs, while accounting for the practical reality of incomplete knowledge. Our contributions are as follows:
@@
  • We present a collection of novel and existing (from Jedwabny et al. (2020)) properties for GS in SGs and discuss their suitability in different settings.
  • Using these properties, we conduct a comprehensive theoretical analysis of our two proposed instantiations, two instantiations of the T-norm semantics Jedwabny et al. (2020) and two existing GS for QBAFs, demonstrating the strengths and limitations of each approach.
@@ 2 Preliminaries
For simplicity and ease of presentation, we restrict attention to statements with conjunctive premises, similarly to Jedwabny et al. (2020). In §3.1, we will define our methodology generally, before instantiating it for these restricted statements, pointing towards other instantiations which could be used for more complex compounds, such as those in conjunctive normal form.
@@
$\bullet$ $\alpha_{1}$ supports $\alpha_{2}$ iff $\exists x\in Prem(\alpha_{2})$ s.t. $\Psi_{1}\equiv x$;
$\bullet$ $\alpha_{1}$ attacks $\alpha_{2}$ iff $\exists x\in Prem(\alpha_{2})$ s.t. $\Psi_{1}\equiv\neg x$.
We can then generate an SG structured with respect to the support and attack relations between statements, as in the works by Hecham et al. Hecham et al. (2018) and Jedwabny et al. Jedwabny et al. (2020). However, differently to these works, we include the quantitative weights of each statement within the SG, rather than applying them separately. This reformulation allows for direct comparisons with properties of QBAFs Baroni et al. (2019), which provided inspiration for our properties.
@@
$\bullet$ $\mathcal{X}$ is a set of statements;
$\bullet$ $\mathcal{A}\!\subseteq\!\mathcal{X}\!\times\!\mathcal{X}$ is s.t. $\forall\alpha_{1},\!\alpha_{2}\!\in\!\mathcal{X}$, $(\alpha_{1},\alpha_{2})\!\in\!\mathcal{A}$ iff $\alpha_{1}$ attacks $\alpha_{2}$;
$\bullet$ $\mathcal{S}\!\subseteq\!\mathcal{X}\!\times\!\mathcal{X}$ is s.t. $\forall\alpha_{1},\!\alpha_{2}\!\in\!\mathcal{X}$, $(\alpha_{1},\!\alpha_{2})\!\in\!\mathcal{S}$ iff $\alpha_{1}$ supports $\alpha_{2}$;
$\bullet$ $\tau:\mathcal{X}\rightarrow[0,1]$ is a weight over the statements.
@@
An SG thus gives a restricted form of structured argumentation framework with conjunctive premises, undercutting attacks, deductive/evidential supports and weights. Note that weights may be fixed, i.e., identical for all arguments which are considered a-priori equal as in classical argumentation frameworks, or variable, representing, e.g., different attributes such as source authority, premise credibility, or goal importance. [^3] While methods for learning these weights represent an important research direction, e.g., as in Rago et al. (2025) within the setting of review aggregation, they fall outside the scope of this work. We envisage a usage of weights similar to that in Alsinet et al. (2008) within the realm of defeasible logic programming García and Simari (2004), where weighted certainties on arguments (including facts) are used to introduce possibilistic argumentation. Also, we focus on acyclic graphs, following several works on GS Rago et al. (2016); de Tarlé et al. (2022), leaving cyclic graphs to future work.
@@
In essence, GS offer a principled approach to evaluating statements quantitatively. We refer to any statement $\alpha\in\mathcal{X}$ with: $\sigma^{\mathcal{G}}(\alpha)=1$ as accepted, and $\sigma^{\mathcal{G}}(\alpha)=0$ as rejected.
Next, we recall two notions from Jedwabny et al. (2020). [^5] We omit the acyclicity condition since we use acyclic graphs.
@@
$\bullet$ $\alpha_{1}\in T$;
$\bullet$ $\forall\alpha_{2}\!\in\!T$, $\forall x\!\in\!Prem(\alpha_{2})$, $\exists\alpha_{3}\!=\!\langle\Phi_{3},\Psi_{3}\rangle\!\in\!T$ s.t. $\Psi_{3}\equiv x$;
$\bullet$ $\nexists\alpha_{2},\alpha_{3}\in T$ s.t. $(\alpha_{2},\alpha_{3})\in\mathcal{A}$;
$\bullet$ $\nexists T^{\prime}\subset T$ s.t. $T^{\prime}$ is a CST for $\alpha_{1}$ (minimal wrt set inclusion).
@@
Then, Jedwabny et al. Jedwabny et al. (2020) introduce the following GS.
@@
In this paper, we instantiate two T-norm semantics. For the first, which we name the T-norm-p semantics, denoted by $\sigma_{T_{p}}$, we let, as in the illustrations by Jedwabny et al. Jedwabny et al. (2020): $\otimes$ be the product, i.e., $\otimes(v_{1},v_{2})=v_{1}\times v_{2}$; $\oplus$ be the probabilistic sum, i.e., $\otimes(v_{1},v_{2})=v_{1}+v_{2}-(v_{1}\times v_{2})$; and $\neg$ be standard negation, i.e., $\neg(v_{1})=1-v_{1}$. For the second, which we name the T-norm-m semantics, denoted by $\sigma_{T_{m}}$, we let: $\otimes$ be the minimum, i.e., $\otimes(v_{1},v_{2})=\min(v_{1},v_{2})$; $\oplus$ be the maximum, i.e., $\otimes(v_{1},v_{2})=\max(v_{1},v_{2})$; and $\neg$ be standard negation as for the T-norm-p semantics.
We refer informally to any GS that takes into account the premises of any statement as a structured GS, and any that does not as an abstract GS. It should be noted that while abstract GS can be applied to SGs, structured GS cannot be applied to frameworks such as QBAFs, since the statements therein are abstract entities.
We now recap two abstract GS from the literature, as defined by Rago et al. Rago et al. (2016) and Potyka Potyka (2018), respectively.
@@
$\bullet$ $\Sigma$ is s.t. $\Sigma(())=0$, where ()()( ) is an empty sequence, and, for $v_{1},\ldots,v_{n}\in[0,1]$ ($n\geq 1$), if $n=1$, then $\Sigma((v_{1}))=v_{1}$; if $n=2$, then $\Sigma((v_{1},v_{2}))=v_{1}+v_{2}-v_{1}\cdot v_{2}$; and if $n>2$, then $\Sigma((v_{1},\ldots,v_{n}))=\Sigma(\Sigma((v_{1},\ldots,v_{n-1})),v_{n})$;
$\bullet$ $c$ is s.t., for $v^{0},v^{-},v^{+}\in[0,1]$, if $v^{-}\geq v^{+}$, then $c(v^{0},v^{-},v^{+})=v^{0}-v^{0}\cdot|v^{+}-v^{-}|$ and if $v^{-}<v^{+}$, then $c(v^{0},v^{-},v^{+})=v^{0}+(1-v^{0})\cdot|v^{+}-v^{-}|$.
@@ 3 Gradual Semantics for Statement Graphs
$\bullet$ complete iff $\forall\alpha_{2}\in\mathcal{X}$, where $\alpha_{2}=\alpha_{1}$ or there exists a path from $\alpha_{2}$ to $\alpha_{1}$ via $\mathcal{S}$, $\mathcal{T}^{\mathcal{G}}(\alpha_{2})\neq\emptyset$;
$\bullet$ partially-complete iff it is not complete and $\mathcal{T}^{\mathcal{G}}(\alpha_{1})\neq\emptyset$;
$\bullet$ incomplete otherwise, i.e., iff $\mathcal{T}^{\mathcal{G}}(\alpha_{1})=\emptyset$.
Note that partially complete and incomplete statements are in the same spirit as potential arguments in ABA Toni (2013), but the latter are used for computational purposes, as steps towards complete statements, rather than as representations in their own right, as we do.
@@
Consider an SG in which $\mathcal{X}=\{\alpha_{1},\alpha_{2},\alpha_{3}\}$, $\alpha_{1}=\langle b,a\rangle$, $\alpha_{2}=\langle c,b\rangle$ and $\alpha_{3}=\langle d,b\rangle$, giving $\mathcal{S}=\{(\alpha_{2},\alpha_{1}),(\alpha_{3},\alpha_{1})\}$ and $\mathcal{A}=\emptyset$ (the $\tau$ values are irrelevant here). In this first case, $\alpha_{1}$ incomplete as it has no CSTs. If we add the statement $\alpha_{4}=\langle\top,c\rangle$, leading to $\mathcal{S}=\{(\alpha_{2},\alpha_{1}),(\alpha_{3},\alpha_{1}),(\alpha_{4},\alpha_{2})\}$, $\alpha_{1}$ becomes partially complete as it has one CST, but $\alpha_{3}$, which has a path to $\alpha_{1}$, does not have a CST. Finally, if we add the statement $\alpha_{5}=\langle\top,d\rangle$, leading to $\mathcal{S}=\{(\alpha_{2},\alpha_{1}),(\alpha_{3},\alpha_{1}),(\alpha_{4},\alpha_{2}),(\alpha_{5},\alpha_{3})\}$, $\alpha_{1}$ becomes complete as all statements with a path to it have CSTs. Crucial here is the fact that the T-norm semantics, in the incomplete case ignores the (potentially useful) information from both $\alpha_{2}$ and $\alpha_{3}$, regardless of $\tau$, while in the partially-complete case it ignores that of $\alpha_{3}$.

### v6

@@ 1 Introduction
Argumentation frameworks have emerged as powerful tools for reasoning about and resolving conflicting information in complex environments (see (??) for overviews). In recent years, gradual semantics (GS) have shown great promise in extending these frameworks, particularly in the realm of quantitative bipolar argumentation frameworks (QBAFs) (?), where arguments have weights and may be related by attack and support relations. These approaches thus allow for both negative and positive, respectively, influences between arguments to be accounted for, arguably aligning more closely with human judgment than traditional argumentation frameworks (?). GS have found applications in diverse (and often human-centric) areas, from judgmental forecasting (?) to explainable AI (??), by allowing for quantitative, and thus in some cases more nuanced, evaluations of arguments beyond the traditional accept/reject dichotomy in extension-based semantics (??). These quantitative evaluations, typically referred to as strengths, can be particularly useful in cases where the information in argumentation frameworks is incomplete, since strengths can account for the resulting uncertainty quantitatively.
@@
However, most of the existing work on GS (see (?) for an overview) has focused on formalisms based on abstract arguments (?), treated as atomic entities without internal structure. This abstraction, while useful in many contexts, fails to capture the rich logical structure that is often present in real-world arguments. Meanwhile, various forms of structured argumentation frameworks (see (?) for an overview) allow to represent the internal composition of arguments and the relationships between their components, and thus offer a more expressive and realistic model in agent-based domains, such as user modeling  (?), scientific debates (?), and model reconciliation settings (?). However, this greater reliance on the logical (e.g., background or contextual) knowledge can lead to more problematic cases where this information is incomplete. Yet, it is only recently that the study of GS for structured argumentation has received attention, e.g., for deductive argumentation (?) by Heyninck et al. (?), for assumption-based argumentation (ABA) (?) by Skiba et al. (?), for ASPIC+ (?) by Spaans (?) and Prakken (?), and for a restricted type of structured argumentation in the form of statement graphs (SGs) (?) by Jedwabny et al. (?).
We selected SGs as our targeted form of structured argumentation in this paper. We chose this starting point for our analysis, rather than other forms of structured argumentation, because they naturally accommodate attack and support relations and their GS are thus naturally relatable to those of QBAFs, which are the most widely studied form of argumentation when it comes to GS (?). This allows us to undertake an interesting analysis of properties of our GS in direct comparison with those for QBAFs, given their widespread application.
For an illustration of the need for GS in structured argumentation, Figure 1 shows a debate about climate change as an SG. SGs use a restricted type of structured arguments in the form of statements, each consisting of a premise (a compound of literals, which may possibly be $\top$, i.e., true) and a claim (a literal). A typical statement ($\alpha_{1}$) may combine evidence about rising CO2 emissions ($a$) and increasing global temperatures ($b$), given that they are both supported by facts ($\alpha_{2}$ and $\alpha_{3}$), to support the claim that climate change is human-caused ($c$). Such an SG may be the result of applying argument mining to text, e.g., as overviewed in  (?). Such argument mining techniques coupled with suitable GS for SGs could allow for better more expressive representation of arguments while handling incomplete information in the modelling of debates, a rapidly growing research area which utilises many forms of argumentation (??????), especially given the groundbreaking potential of large language models (?).
@@
The GS proposed by Jedwabny et al. (?) for SGs uses T-norms and T-conorms from fuzzy logic to aggregate the strengths of a statement’s components, but this GS faces challenges when dealing with incomplete information since it relies on an assumption of complete knowledge and presence of a statement’s supporting elements. This requirement can limit its applicability in real-world scenarios, where such information is almost always incomplete. For example, in Figure 1, the strength of the statement $\alpha_{1}$ is reduced (e.g., in the eyes of a modelled audience) by a statement ($\alpha_{4}$) questioning the reliability of CO2 measurements, without being evidenced (there is no support for $d$). To date, existing approaches to GS in structured argumentation struggle to handle such incompleteness, potentially leading to oversimplified or incorrect conclusions, and thus poor modelling of debates.
To address this limitation, we introduce a novel methodology for GS in SGs that can naturally accommodate complete and incomplete information. Our approach is modular in that it allows for different GS to be instantiated based on the requirements of a given task and context. Specifically, our GS can be instantiated with any existing GS for QBAFs. This existing GS is used to dialectically evaluate the literals in a premise, based on their attacking and supporting evidence, before these “intermediate” evaluations are aggregated (depending on the statement’s construction) to give a final strength for the statement.
Thus, our methodology combines some of the expressiveness of structured argumentation with the flexibility of GS for QBAFs, while accounting for the practical reality of incomplete knowledge. Our contributions are as follows:
@@
  • We present a collection of novel and existing (by Jedwabny et al. (?)) properties for GS in SGs and discuss their suitability in different settings.
  • Using these properties, we conduct a comprehensive theoretical analysis of our two proposed instantiations, two instantiations of the T-norm semantics (?) and two existing GS for QBAFs, demonstrating the strengths and limitations of each approach.
2 Related Work
GS have received significant attention within abstract argumentation in recent years, given their professed alignment with human reasoning (??) and applicability to real-world settings (??). This has given rise to a rich field of research, with a broad repertoire of GS (???), in-depth analyses of GS’ behaviour (????), and a number of open-source computational toolsets (??). One issue in this area is that many abstract GS, given their recursive nature, do not satisfy the uniqueness property in assigning strengths in cyclic graphs (??). This must be considered if our acyclicity restriction is relaxed in future work.
With regards to other approaches to GS in structured argumentation, our approach differs from existing work as follows. Within ASPIC+, Spaans (?) introduced initial strength functions, with corresponding properties, though they use weights on rules and do not consider explicit support relations. Further, Prakken (?) gives a formal model of argument strength and deploys it. However, the main difference between this work and our approach is our leveraging of existing bipolar abstract GS for handling incomplete information. We leave to future work the assessment of this approach with respect to our proposed properties. Also for ASPIC+, extension-based semantics have been used for cases with incomplete information (?).
Meanwhile, Heyninck et al. (?) compare properties of GS with those concerning culpability when applied to abstract argumentation frameworks extracted from deductive argumentation, whereas we lift abstract GS to the structured level. Skiba et al. (?) assess a family of GS for ranking arguments in structured argumentation, specifically ABA. Both of these works do not consider the support relation. Finally, Amgoud and Ben-Naim (?) give a method for ranking logic-based argumentation frameworks, but do not consider an explicit strength or a support relation.
Finally, a related but distinct subfield is probabilistic argumentation (???????), but in all of these works, either support relations are not considered or strengths are associated with probabilities and/or are not assigned to individual arguments.
@@ 3 Preliminaries
For simplicity and ease of presentation, we restrict attention to statements with conjunctive premises, similarly to (?). In §4.1, we will define our methodology generally, before instantiating it for these restricted statements, pointing towards other instantiations which could be used for more complex compounds, such as those in conjunctive normal form.
@@
  • $\alpha_{1}$ supports $\alpha_{2}$ iff $\exists x\in Prem(\alpha_{2})$ s.t. $\Psi_{1}\equiv x$;
  • $\alpha_{1}$ attacks $\alpha_{2}$ iff $\exists x\in Prem(\alpha_{2})$ s.t. $\Psi_{1}\equiv\neg x$.
We can then generate an SG structured with respect to the support and attack relations between statements, as in the works by Hecham et al. (?) and Jedwabny et al. (?). [^2] Note that we chose to use SGs similarly to Hecham et al. (?) and Jedwabny et al. (?) in order to allow for direct comparisons with the gradual semantics in the latter work. Their focus on attacks and supports on premises, rather than conclusions, also aligns with other works in the literature, e.g., ABA. However, differently to these works, we include the quantitative weights of each statement within the SG, rather than applying them separately. This reformulation allows for direct comparisons with properties of QBAFs (?), which provided inspiration for some of our properties.
@@
  • $\mathcal{X}$ is a set of statements;
  • $\mathcal{A}\!\subseteq\!\mathcal{X}\!\times\!\mathcal{X}$ is s.t. $\forall\alpha_{1},\!\alpha_{2}\!\in\!\mathcal{X}$, $(\alpha_{1},\alpha_{2})\!\in\!\mathcal{A}$ iff $\alpha_{1}$ attacks $\alpha_{2}$;
  • $\mathcal{S}\!\subseteq\!\mathcal{X}\!\times\!\mathcal{X}$ is s.t. $\forall\alpha_{1},\!\alpha_{2}\!\in\!\mathcal{X}$, $(\alpha_{1},\!\alpha_{2})\!\in\!\mathcal{S}$ iff $\alpha_{1}$ supports $\alpha_{2}$;
  • $\tau:\mathcal{X}\rightarrow[0,1]$ is a weight over the statements.
@@
An SG thus gives a restricted form of structured argumentation framework with conjunctive premises, undercutting attacks, deductive/evidential supports and weights. Note that weights may be fixed, i.e., identical for all arguments which are considered a-priori equal as in classical argumentation frameworks, or variable, representing, e.g., different attributes such as source authority, premise credibility, or goal importance. [^4] While methods for learning these weights represent an important research direction, e.g., as in (?) within the setting of review aggregation, they fall outside the scope of this work. We envisage a usage of weights similar to that in (?) within the realm of defeasible logic programming (?), where weighted certainties on arguments (including facts) are used to introduce possibilistic argumentation. Also, we focus on acyclic graphs, following several works on GS (??), leaving cyclic graphs to future work.
@@
In essence, GS offer a principled approach to evaluating statements quantitatively. We refer informally to any GS that takes into account the premises of any statement as a structured GS, and any that does not as an abstract GS. It should be noted that while abstract GS can be applied to SGs, structured GS cannot be applied to frameworks using abstract argumentation such as QBAFs, since the statements therein are abstract entities.
We refer to any statement $\alpha\in\mathcal{X}$ with $\sigma^{\mathcal{G}}(\alpha)=1$ as having/being top-strength, and $\sigma^{\mathcal{G}}(\alpha)=0$ as having/being bottom-strength. We note here that other notions of acceptance and rejection, e.g., with thresholds over the strength range as in (??), would be possible in our setting, but we leave the study of this to future work.
Next, we recall two notions from (?). [^6] We omit the acyclicity condition since we use acyclic graphs.
@@
  • $\alpha_{1}\in T$;
  • $\forall\alpha_{2}\!\in\!T$, $\forall x\!\in\!Prem(\alpha_{2})$, $\exists\alpha_{3}\!=\!\langle\Phi_{3},\Psi_{3}\rangle\!\in\!T$ s.t. $\Psi_{3}\equiv x$;
  • $\nexists\alpha_{2},\alpha_{3}\in T$ s.t. $(\alpha_{2},\alpha_{3})\in\mathcal{A}$;
  • $\nexists T^{\prime}\subset T$ s.t. $T^{\prime}$ is a CST for $\alpha_{1}$ (minimal wrt set inclusion).
@@
Then, Jedwabny et al. (?) introduce the following GS.
@@
While the T-norm semantics is modular wrt the De Morgan triple, the methodology is based on CSTs, and thus it can be seen that it inherently relies on complete information.
In this paper, we instantiate two T-norm semantics. For the first, which we name the T-norm-p semantics, denoted by $\sigma_{T_{p}}$, we let, as in the illustrations by Jedwabny et al. (?): $\otimes$ be the product, i.e., $\otimes(v_{1},v_{2})=v_{1}\times v_{2}$; $\oplus$ be the probabilistic sum, i.e., $\oplus(v_{1},v_{2})=v_{1}+v_{2}-(v_{1}\times v_{2})$; and $\neg$ be standard negation, i.e., $\neg(v_{1})=1-v_{1}$. For the second, which we name the T-norm-m semantics, denoted by $\sigma_{T_{m}}$, we let: $\otimes$ be the minimum, i.e., $\otimes(v_{1},v_{2})=\min(v_{1},v_{2})$; $\oplus$ be the maximum, i.e., $\oplus(v_{1},v_{2})=\max(v_{1},v_{2})$; and $\neg$ be standard negation as for the T-norm-p semantics.
We now recap two abstract GS from the literature, as defined by Rago et al. (?) and Potyka (?), respectively.
@@
  • $\Sigma$ is s.t. $\Sigma(())=0$, where ()()( ) is an empty sequence, and, for $v_{1},\ldots,v_{n}\in[0,1]$ ($n\geq 1$), if $n=1$, then $\Sigma((v_{1}))=v_{1}$; if $n=2$, then $\Sigma((v_{1},v_{2}))=v_{1}+v_{2}-v_{1}\cdot v_{2}$; and if $n>2$, then $\Sigma((v_{1},\ldots,v_{n}))=\Sigma(\Sigma((v_{1},\ldots,v_{n-1})),v_{n})$;
  • $c$ is s.t., for $v^{0},v^{-},v^{+}\in[0,1]$, if $v^{-}\geq v^{+}$, then $c(v^{0},v^{-},v^{+})=v^{0}-v^{0}\cdot|v^{+}-v^{-}|$ and if $v^{-}<v^{+}$, then $c(v^{0},v^{-},v^{+})=v^{0}+(1-v^{0})\cdot|v^{+}-v^{-}|$.
@@
While these abstract GS do not take into account the structure of statements, they do consider a weight on each statement, which gives an intrinsic strength which may be used as a starting point for calculating statement strengths, e.g., in the case of incomplete information.
@@ 4 Gradual Semantics for Statement Graphs
  • complete iff $\forall\alpha_{2}\in\mathcal{X}$, where $\alpha_{2}=\alpha_{1}$ or there exists a path from $\alpha_{2}$ to $\alpha_{1}$ via $\mathcal{S}$, $\mathcal{T}^{\mathcal{G}}(\alpha_{2})\neq\emptyset$;
  • partially-complete iff it is not complete and $\mathcal{T}^{\mathcal{G}}(\alpha_{1})\neq\emptyset$;
  • incomplete otherwise, i.e., iff $\mathcal{T}^{\mathcal{G}}(\alpha_{1})=\emptyset$.
Note that partially-complete and incomplete statements are in the same spirit as potential arguments in ABA (?), but the latter are used for computational purposes, as steps towards complete statements, rather than as representations in their own right, as we do.
@@
Consider an SG, as illustrated in Figure 2, in which $\mathcal{X}=\{\alpha_{1},\alpha_{2},\alpha_{3}\}$, $\alpha_{1}=\langle b,a\rangle$, $\alpha_{2}=\langle c,b\rangle$ and $\alpha_{3}=\langle d,b\rangle$, giving $\mathcal{S}=\{(\alpha_{2},\alpha_{1}),(\alpha_{3},\alpha_{1})\}$ and $\mathcal{A}=\emptyset$ (the $\tau$ values are irrelevant here). In this first case, $\alpha_{1}$ incomplete as it has no CSTs. If we add the statement $\alpha_{4}=\langle\top,c\rangle$, leading to $\mathcal{S}=\{(\alpha_{2},\alpha_{1}),(\alpha_{3},\alpha_{1}),(\alpha_{4},\alpha_{2})\}$, $\alpha_{1}$ becomes partially-complete as it has one CST, but $\alpha_{3}$, which has a path to $\alpha_{1}$, does not have a CST. Finally, if we add the statement $\alpha_{5}=\langle\top,d\rangle$, leading to $\mathcal{S}=\{(\alpha_{2},\alpha_{1}),(\alpha_{3},\alpha_{1}),(\alpha_{4},\alpha_{2}),(\alpha_{5},\alpha_{3})\}$, $\alpha_{1}$ becomes complete as all statements with a path to it have CSTs. Crucial here is the fact that in the incomplete case the T-norm semantics ignores the (potentially useful) information from both $\alpha_{2}$ and $\alpha_{3}$, regardless of $\tau$, while in the partially-complete case it ignores that of $\alpha_{3}$.
Figure 2: SG from Example 1 with initial (added) statements shown in solid (striped, respectively) turquoise, no attacks, initial (added) supports as solid (dashed, respectively) green arrows and some arbitrary $\tau$ (not shown here). The statement $\alpha_{1}\in\mathcal{X}$ is incomplete when $\mathcal{X}=\{\alpha_{1},\alpha_{2},\alpha_{3}\}$, but is partially-complete when $\alpha_{4}$ (or $\alpha_{5}$) is added and complete when both $\alpha_{4}$ and $\alpha_{5}$ are added.